# Quantitative surgery and total mean curvature

### v1

@@ 5. Surgery bending in arbitrary backgrounds
By the Gauß-Equation we have
(5.9) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\underbrace{\mathrm{scal}(M\times[0,1],g+\mathrm{d}t^{2})}_{=\mathrm{scal}(M,g)}\\ &\qquad={}H^{2}(\Sigma_{\varrho},g+\mathrm{d}t^{2})-|A(\Sigma_{\varrho},g+\mathrm{d}t^{2})|^{2}-2\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{\varrho},\nu_{\varrho}),\end{split}$
where $\nu_{\varrho}$ is the upwards unit normal vector of $\Sigma_{\varrho}$ with respect to the ambient metric $g+\mathrm{d}t^{2}$. In order to estimate the summands from the right-hand side of (5.9), consider the symmetric $2$-tensor $\gamma$ on $S$ defined as the first order Taylor expansion of $g$ in normal direction, that is
(5.10) $\gamma_{\alpha\beta}=g_{\alpha\beta}|_{S}+\sum_{i=1}^{n-k}x^{i}g_{\alpha\beta,i}|_{S}$
which we extend canonically to $B_{R}^{n-k}$. Since the tensor $\gamma$ precisely describes the $1$-jet of the metric $g$ along $S$, there exists an $\varepsilon>0$ such that the restriction of $\gamma$ to the disk-bundle of radius $\varepsilon$ is positive definite and hence a Riemannian metric. If $r$ denotes the radial component of the fiber of the tubular neighbourhood $D(S)$ in $M$, we have
@@
since $g_{\alpha\beta}(x)=g_{\alpha\beta}(p)+x^{k}g_{\alpha\beta,k}(p)+O(r^{2})$ by the Taylor remainder estimate. The advantage of working with the background metric $\gamma+\mathrm{d}t^{2}$ is that the expressions for $H$ and $|A|$ can be readily compared to the corresponding expressions for Euclidean handles as described in Section 4.
@@
$|\nu_{g}-\nu_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\lesssim{}$ $r^{2},\qquad$ $|A_{g}-A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\lesssim{}$ $(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot r^{2},$
where the latter identity follows from $A_{g}(e_{i},e_{j})=-\langle\nabla^{g+dt^{2}}_{e_{i}}e_{j},\nu_{g}\rangle_{g}$ and $|\Gamma^{k}_{ij}|_{g}\lesssim r^{2}$ for the Christoffel symbols $\Gamma_{ij}^{k}=\tfrac{1}{2}g^{kl}(\nabla^{\gamma}_{i}g_{lj}+\nabla^{\gamma}_{j}g_{li}+\nabla^{\gamma}_{l}g_{ij})$.
@@
$={}$ $\bigl|(|A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})(|A_{g}|_{g+\mathrm{d}t^{2}}+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\bigr|$
$={}$ $\bigl||A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr|^{2}+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigl||A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr|$
$\leq{}$ $\varrho^{2}(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho$
$\leq{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\left(\varrho(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\right)$
@@
$|H_{g}^{2}-H_{\gamma}^{2}|={}$ $\bigl|(H_{g}-H_{\gamma})(H_{g}+H_{\gamma})\bigr|\leq|H_{g}-H_{\gamma}|^{2}+2|H_{\gamma}|\cdot|H_{g}-H_{\gamma}|$
$\leq{}$ $(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}\cdot\varrho^{2}+n|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho$
$={}$ $(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho((1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho+n|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})$
$\lesssim{}$ $\varrho(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}.$
@@
$\qquad\leq{}\underbrace{\bigl|\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}\bigr|}_{\leq\|\nu_{g}-\nu_{\gamma}\|_{g+\mathrm{d}t^{2}}\lesssim\varrho}+\underbrace{\bigl|\langle\nu_{\gamma},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\bigr|}_{\leq\|g-\gamma\|_{\gamma}\lesssim\varrho}\lesssim{}\varrho$
@@
$\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{g},\nu_{g})={}$ $\|\nu_{g}^{\intercal}\|^{2}\cdot\mathrm{Ric}(M,g)\left(\frac{\nu_{g}^{\intercal}}{\|\nu_{\gamma}^{\intercal}\|},\frac{\nu_{\gamma}^{\intercal}}{\|\nu_{\gamma}^{\intercal}\|}\right)$
$\lesssim{}$ $\|\nu_{g}^{\intercal}\|^{2}\lesssim\|\nu_{\gamma}^{\intercal}\|$
@@
As mentioned above, the advantage of working with the background metric $\gamma+\mathrm{d}t^{2}$ is that the expressions for $H$ and $|A|$ can be readily compared to the corresponding expressions for Euclidean handles as described in Section 4. To make this precise, we first observe that the projection
$\pi:D(S)\times[0,1]\longrightarrow B_{R}^{n-k}(0)\times[0,1]\subseteq\mathbb{R}^{n-k}\times[0,1]$
@@
If we fix a point $x\in\Sigma_{\varrho}$ and choose an adapted $\gamma$-orthonormal frame $\{e_{i},e_{a}\}$ of $\Sigma_{\varrho}$ at $x$, where $\{e_{i}\}$ are normal to $S$ and $\{e_{a}\}$ tangent to $S$, we see that $\{D\pi(e_{i})\}$ form normal coordinates at $\pi(x)\in\overline{\Sigma}_{\varrho}$. By definition of the second fundamental form,
$A_{\gamma}(e_{i},e_{j})={}$ $-\langle\nabla^{\gamma+\mathrm{d}t^{2}}_{e_{i}}e_{j},\nu_{\gamma}\rangle=-\langle\nabla^{\delta+\mathrm{d}t^{2}}_{e_{i}}e_{j},\nu_{\delta}\rangle=A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})(e_{i},e_{j}).$
@@
$|A_{\gamma}(e_{i},e_{a})|\lesssim\|\nu_{\delta}^{\intercal}\|\quad|A_{\gamma}(e_{a},e_{b})|\lesssim\|\nu_{\delta}^{\intercal}\|\quad\text{and}\quad A_{\gamma}(e_{i},e_{j})=A(\overline{\Sigma}_{\varrho},\delta+\mathrm{d}t^{2})(e_{i},e_{j}).$
@@
$|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}\lesssim{}$ $\|\nu_{\delta}^{\intercal}\|^{2}$
@@
$|H_{\gamma}^{2}-H_{\delta}^{2}|={}$ $2|H_{\delta}|\cdot|H_{\gamma}-H_{\delta}|+(H_{\gamma}-H_{\delta})^{2}$
@@
where we used $|H_{\delta}|\leq n|A_{\delta}|$}. Combining these, we get
@@
and therefore $|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}\lesssim\|\nu_{\varrho}^{\intercal}\|\leq 1$. Note that $H_{\delta}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}$ is precisely the scalar curvature of $(\Sigma_{\varrho},g_{\varrho,\delta})$, where $g_{\varrho,\delta}$ is induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$. Thus, we obtain
@@
$\qquad=(H_{g}^{2}-|A_{g}|_{g+\mathrm{d}t^{2}}^{2})-H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}-H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}+H_{\delta}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}$
$\qquad\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}$
$\qquad\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}}+\|\nu_{\delta}^{\intercal}\|)^{2}$
$\qquad\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}$
Putting all the pieces together, we therefore obtain the following inequality, where the constant $C$ might change from line to line, independently of $\varrho$:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g)-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})$
$\qquad=\underbrace{H^{2}_{g}-|A_{g}|^{2}-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})}_{\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}}-2\underbrace{\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{\varrho},\nu_{\varrho})}_{\lesssim\|\nu_{\delta}^{\intercal}\|}$
$\qquad\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}}+\|\nu_{\delta}^{\intercal}\|)^{2}-\|\nu_{\delta}^{\intercal}\|$
$\qquad\gtrsim-(\varrho+\|\nu_{\delta}^{\intercal}\|)\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}$
@@
Next, we compare $\Sigma_{\varrho}$ for different values of $\varrho$. Lemma 4.4 immediately yields:
@@
For every $\mu$ and every $\varrho\in[\varrho_{0},\varrho_{1}]$ there exists diffeomorphisms $\Phi_{\varrho}:\Sigma_{\varrho_{1}}\to\Sigma_{\varrho}$ such that $\Phi_{\varrho_{1}}=\operatorname{\mathrm{id}}_{\Sigma_{\varrho_{1}}}$ and for $g_{\mu,\varrho}$ the metric on $\Sigma_{\varrho}$ induced by $g_{\mu}$ (see Theorem 5.1) the assignment
$\varrho\mapsto e^{\varrho}\cdot\Phi_{\varrho}^{\ast}(g_{\varrho,\mu})$
@@
Finally, we deform the metric $e^{\varrho_{1}}\cdot g_{\mu,\varrho_{1}}$ to a metric which is independent of $\mu$ by linear interpolation of conformal factors:
@@
For $\varrho_{1}>0$ and every $\mu\in(0,1]$, define the path $g_{s}$, $s\in[0,1]$ of metrics on $\Sigma_{\varrho_{1}}$ to be induced by the family
$s\mapsto e^{\varrho_{1}}((1-s)\cdot u+2s)^{\frac{4}{n-2}}\cdot(g+\mathrm{d}t^{2})$
of metrics on $M\times[0,1]$, where $u=(1+t\mu\eta(3t-1))$ as in Theorem 5.1. Then, $g_{s}$ is strictly monotonically increasing in $s$ and satisfies
$\mathrm{scal}(\Sigma_{\varrho_{1}},g_{s})\geq-e^{-\varrho_{1}}(|\mathrm{scal}(M,g)|+\varrho_{1}^{1/4}).$

### v2

@@ 6. Surgery bending in arbitrary backgrounds
By the Gauss equation we have
(6.18) $\begin{split}&\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\underbrace{\mathrm{scal}(M\times[0,1],g+\mathrm{d}t^{2})}_{=\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}}\\ &\qquad={}H(\Sigma_{\varrho},g+\mathrm{d}t^{2})^{2}-|A(\Sigma_{\varrho},g+\mathrm{d}t^{2})|^{2}-2\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{\varrho},\nu_{\varrho}),\end{split}$
where $\nu_{\varrho}$ is the upwards unit normal vector of $\Sigma_{\varrho}$ with respect to the ambient metric $g+\mathrm{d}t^{2}$ and $\mathrm{pr}_{M}\colon M\times[0,1]\to M$ is the projection. Again, we will omit $\circ\mathrm{pr}_{M}$ from now on. In order to estimate the summands from the right-hand side of (6.18), consider the symmetric $2$-tensor $\gamma$ on $\mathcal{S}$ defined as the first order Taylor expansion of $g$ in normal direction, that is
(6.19) $\gamma_{\alpha\beta}=g_{\alpha\beta}|_{\mathcal{S}}+\sum_{i=1}^{k}x^{i}g_{\alpha\beta,i}|_{\mathcal{S}}$
which we extend canonically to $B_{R}^{k}$. Since the tensor $\gamma$ precisely describes the $1$-jet of the metric $g$ along $\mathcal{S}$, there exists an $\varepsilon>0$ such that the restriction of $\gamma$ to the disk-bundle of radius $\varepsilon$ is positive definite and hence a Riemannian metric. If $r$ denotes the radial component of the fiber of the tubular neighbourhood $D(\mathcal{S})$ in $M$, we have
@@
since $g_{\alpha\beta}(x)=g_{\alpha\beta}(p)+x^{i}g_{\alpha\beta,i}(p)+O(r^{2})$ by the Taylor remainder estimate. The advantage of working with the background metric $\gamma+\mathrm{d}t^{2}$ is that the expressions for $H$ and $|A|$ can be readily compared to the corresponding expressions for Euclidean handles as described in Section 5.
@@
$\|\nu_{g}-\nu_{\gamma}\|_{\gamma+\mathrm{d}t^{2}}\lesssim{}$ $r^{2},\qquad$ $|A_{g}-A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\lesssim{}$ $(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot r^{2},$
where the latter identity follows from $A_{g}(e_{i},e_{j})=-\langle\nabla^{g+dt^{2}}_{e_{i}}e_{j},\nu_{g}\rangle_{g}$ and $|\Gamma^{m}_{ij}|_{g}\lesssim r^{2}$ for the Christoffel symbols $\Gamma_{ij}^{m}=\tfrac{1}{2}g^{ml}(\nabla^{\gamma}_{i}g_{lj}+\nabla^{\gamma}_{j}g_{li}+\nabla^{\gamma}_{l}g_{ij})$.
@@
$\leq{}$ $\bigl||A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr|^{2}+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigl||A_{g}|_{g+\mathrm{d}t^{2}}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr|$
$\lesssim{}$ $\varrho^{2}\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho$
$\lesssim{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\left(\varrho(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})+2|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\right)$
@@
$|H_{g}^{2}-H_{\gamma}^{2}|\leq{}$ $|H_{g}-H_{\gamma}|^{2}+2|H_{\gamma}|\cdot|H_{g}-H_{\gamma}|$
$\lesssim{}$ $\varrho^{2}\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}+n|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho$
$\lesssim{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\bigl((1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})\cdot\varrho+n|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\bigr)$
$\lesssim{}$ $\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}.$
@@
$\qquad\leq{}\underbrace{\bigl|\langle\nu_{g},\partial_{t}\rangle_{g+\mathrm{d}t^{2}}-\langle\nu_{g},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\bigr|}_{\lesssim\|g-\gamma\|_{\gamma}\lesssim\varrho}+\underbrace{\bigl|\langle\nu_{g},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}-\langle\nu_{\gamma},\partial_{t}\rangle_{\gamma+\mathrm{d}t^{2}}\bigr|}_{\lesssim\|\nu_{g}-\nu_{\gamma}\|_{\gamma+\mathrm{d}t^{2}}\lesssim\varrho}\lesssim{}\varrho$
@@
$\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{g},\nu_{g})={}$ $\|\nu_{g}^{\intercal}\|_{g}^{2}\cdot\mathrm{Ric}(M,g_{M})\left(\frac{\nu_{g}^{\intercal}}{\|\nu_{g}^{\intercal}\|_{g}},\frac{\nu_{g}^{\intercal}}{\|\nu_{g}^{\intercal}\|_{g}}\right)$
$\lesssim{}$ $\|\nu_{g}^{\intercal}\|_{g}^{2}\lesssim\|\nu_{\gamma}^{\intercal}\|_{\gamma}+\varrho$
@@
As mentioned above, the advantage of working with the background metric $\gamma+\mathrm{d}t^{2}$ is that the expressions for $H$ and $|A|$ can be readily compared to the corresponding expressions for Euclidean handles from Section 5. To make this precise, we first observe that the projection
$\pi:B_{R}(\mathcal{S})\times[0,1]\longrightarrow B_{R}^{k}(0)\times[0,1]\subseteq\mathbb{R}^{k}\times[0,1]$
@@
If we fix a point $x\in\Sigma_{\varrho}$ and choose an adapted $\gamma$-orthonormal frame $\{e_{i},e_{a}\}$ of $\Sigma_{\varrho}$ at $x$, where $\{e_{i}\}$ are normal to $\mathcal{S}$ and $\{e_{a}\}$ tangent to $\mathcal{S}$, we see that $\{D\pi(e_{i})\}$ form normal coordinates at $\pi(x)\in\overline{\Sigma}_{\varrho}$. By definition of the second fundamental form,
$A_{\gamma}(e_{i},e_{j})={}$ $-\langle\nabla^{\gamma+\mathrm{d}t^{2}}_{e_{i}}e_{j},\nu_{\gamma}\rangle=-\langle\nabla^{\delta+\mathrm{d}t^{2}}_{e_{i}}e_{j},\nu_{\delta}\rangle=A_{\delta}(e_{i},e_{j}).$
@@
$|A_{\gamma}(e_{i},e_{a})|\lesssim\|\nu_{\delta}^{\intercal}\|\quad|A_{\gamma}(e_{a},e_{b})|\lesssim\|\nu_{\delta}^{\intercal}\|\quad\text{and}\quad A_{\gamma}(e_{i},e_{j})=A_{\delta}(e_{i},e_{j}).$
@@
$\bigl||A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}\bigr|\lesssim{}$ $\|\nu_{\delta}^{\intercal}\|^{2}\leq\|\nu_{\delta}^{\intercal}\|$
@@
$|H_{\gamma}^{2}-H_{\delta}^{2}|={}$ $|2H_{\delta}\cdot(H_{\gamma}-H_{\delta})+(H_{\gamma}-H_{\delta})^{2}|$
@@
where we used ${|H_{\delta}|\leq n|A_{\delta}|}\lesssim|A_{\delta}|$. Combining these, we get
@@
and therefore $|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}\lesssim\|\nu_{\varrho}^{\intercal}\|\leq 1$ and $1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}\lesssim 1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}}$. Note that $H_{\delta}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2}$ is precisely the scalar curvature of $(\Sigma_{\varrho},g_{\varrho,\delta})$, where $g_{\varrho,\delta}$ is induced by the Euclidean metric $\delta+\mathrm{d}t^{2}$. Thus, we obtain
@@
$\quad=(H_{g}^{2}-|A_{g}|_{g+\mathrm{d}t^{2}}^{2})-(H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2})+(H_{\gamma}^{2}-|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}}^{2})-(H_{\delta}^{2}-|A_{\delta}|_{\delta+\mathrm{d}t^{2}}^{2})$
$\quad\gtrsim-\varrho\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})^{2}-\|\nu_{\delta}^{\intercal}\|\cdot(1+|A_{\gamma}|_{\gamma+\mathrm{d}t^{2}})$
$\quad\gtrsim-\varrho\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}-\|\nu_{\delta}^{\intercal}\|\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})$
Putting all the pieces together, we therefore obtain:
$\mathrm{scal}(\Sigma_{\varrho},g_{\varrho})-\mathrm{scal}(M,g_{M})-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})$
$\qquad=H^{2}_{g}-|A_{g}|^{2}-\mathrm{scal}(\overline{\Sigma}_{\varrho},g_{\varrho,\delta})-2\underbrace{\mathrm{Ric}(M\times[0,1],g+\mathrm{d}t^{2})(\nu_{\varrho},\nu_{\varrho})}_{\lesssim\|\nu_{\delta}^{\intercal}\|+\varrho}$
$\qquad\gtrsim-\varrho\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}-\|\nu_{\delta}^{\intercal}\|\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})-\|\nu_{\delta}^{\intercal}\|+\varrho$
$\qquad\gtrsim-\varrho\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})^{2}-\|\nu_{\delta}^{\intercal}\|\cdot(1+|A_{\delta}|_{\delta+\mathrm{d}t^{2}})$
@@
Next, we compare $\Sigma_{\varrho}$ for different values of $\varrho$.
@@
For every $\kappa\in(0,\varrho_{1}]$ there exists a smooth path of diffeomorphisms $\Phi_{\varrho}:\Sigma_{\varrho_{1}}\to\Sigma_{\varrho}$, $\varrho_{0}\leq\varrho\leq\varrho_{1}$, such that $\Phi_{\varrho_{1}}=\operatorname{\mathrm{id}}_{\Sigma_{\varrho_{1}}}$ and for $g_{\kappa,\varrho}$ the metric on $\Sigma_{\varrho}$ induced by $g_{\kappa}$ (see Theorem 6.1) the assignment
$\varrho\mapsto e^{\varrho}\cdot\Phi_{\varrho}^{\ast}(g_{\varrho,\kappa})$
@@
Proof.
Recall from (5.5) the family of reparametrizations
$\varphi_{\varrho_{1},\varrho}(\tau)=\tau-q_{\varrho_{1},\varrho}\int_{0}^{\tau}\beta\left(s-\frac{1}{2}\right)ds$
and their lift to $\overline{\Sigma}_{\varrho}$ and $\Sigma_{\varrho}$ which, by slight abuse of notation, we both denote by $\Phi_{\varrho_{1},\varrho}$. First, we show that the proof in Proposition 5.4 still goes through in case the Euclidean background metric $\delta+\mathrm{d}t^{2}$ is replaced with $g+dt^{2}$ restricted to the fiber of the tubular neighbourhood $B_{R}(\mathcal{S})\times[0,1]$. We denote this restriction of $g_{M}+dt^{2}$ by $\hat{g}$. As in the proof of Proposition 5.4 we have
$\Phi_{\varrho_{1},\varrho}^{\ast}g_{\overline{\Sigma}_{\varrho},\hat{g}}(x,\tau)=\hat{g}(x,\varphi_{\varrho_{1},\varrho}(\tau))\bigl(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau),\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau)\bigr)\mathrm{d}\tau^{2}+g_{S^{k-1}_{r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))},\hat{g}}$.
Here $g_{S^{k-1}_{r},\hat{g}}$ is the metric induced by $\hat{g}$ on a sphere of radius $r$ where the radius is measured with respect to $\delta$. For $\varrho_{1}$ sufficiently small, the metric on the spheres $S^{k}_{r}$ is increasing as long as the radius $r\leq\varrho_{1}$ increases. Consequently,
$\partial_{\varrho}g_{S^{k-1}_{r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))},\hat{g}}\geq 0.$
Moreover, we have $|\partial_{\varrho}\varphi_{\varrho_{1},\varrho}(\tau)|\lesssim\varrho^{1/4}\ll 1$. Since the derivatives of $\hat{g}$ are bounded, this yields as in the proof of Proposition 5.4
$\partial_{\varrho}\Bigl(\hat{g}(x,\varphi_{\varrho_{1},\varrho}(\tau))\bigl(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau),\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau)\bigr)\mathrm{d}\tau^{2}\Bigr)$
$\qquad\geq-\frac{1}{2}\hat{g}(x,\varphi_{\varrho_{1},\varrho}(\tau))(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau),\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau))\mathrm{d}\tau^{2}.$
Consequently,
$\partial_{\varrho}\left[e^{\varrho/2}\left(\hat{g}(x,\varphi_{\varrho_{1},\varrho}(\tau))(\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau),\partial_{\tau}\varphi_{\varrho_{1},\varrho}(\tau))\mathrm{d}\tau^{2}+g_{S^{k-1}_{r_{\varrho}(\varphi_{\varrho_{1},\varrho}(\tau))},\hat{g}}\right)\right]\geq 0$
and Proposition 5.4 still holds for the background metric $\hat{g}$. Next, we note that the other components of the metric $g_{\varrho}$ are independent of $\varrho$ except for the basepoint change caused by the reparametrization $\Phi$. Since $|\partial_{\varrho}\varphi_{\varrho_{1},\varrho}(\tau)|\lesssim\varrho^{1/4}$, and since the geometry of both $M$ and $\mathcal{S}$ are bounded, the corresponding $\varrho$-derivatives of these remaining components are negligible small. Moreover, recall that the conformal factor $u_{\kappa}(x,t)=1+t\tilde{\kappa}\eta\left(\alpha^{-1}t-1\right)\leq 1+\varrho_{1}$ is fixed and again that $|\partial_{\varrho}\varphi_{\varrho_{1},\varrho}(\tau)|\ll 1$. Hence, the result follows. Finally, we remark that the above estimates still hold in view of the smoothing construction performed at the end of Section 5. ∎
Finally, we deform the metric $e^{\varrho_{1}}\cdot g_{\kappa,\varrho_{1}}$ to a metric which is independent of $\kappa$ by linear interpolation of conformal factors:
@@
For $\varrho_{1}>0$ and every $\kappa\in(0,\varrho_{1}]$, define the path $g_{s}$, $s\in[0,1]$ of metrics on $\Sigma_{\varrho_{1}}$ to be induced by the family
$s\mapsto e^{\varrho_{1}}\bigl((1-s)\cdot u+2s\bigr)^{\frac{4}{n-2}}\cdot g_{\varrho_{1}}$
of metrics on $M\times[0,1]$, where $u=(1+t\tilde{\kappa}\eta(\alpha^{-1}t-1))$ as in Theorem 6.1. Then, $g_{s}$ is strictly monotonically increasing in $s$ and satisfies
$\mathrm{scal}(\Sigma_{\varrho_{1}},g_{s})\geq-(|\mathrm{scal}(M,g_{M})\circ\mathrm{pr}_{M}|+\varrho_{1}^{1/4}).$